OpenStudy (anonymous):

The product of two consecutive integers is 48 more than the larger. Find the integers.

OpenStudy (anonymous):

@zzr0ck3r would the smallest consecutive even integer be x and the largest be x+4?

OpenStudy (welshfella):

x(x + 1) = x +1 + 48

OpenStudy (welshfella):

x^2 + x - x = 49 x^2 = 49

OpenStudy (welshfella):

so x = square root of 49

OpenStudy (zzr0ck3r):

This time we had nothing to do with odds, so two consecutive numbers are x and x+1 like 2,3 4,5 7,8 -2,-3